Helium-neon laser light of wavelength 632.8 nm is sent through a 0.300mm wide single slit. What is the width of central maximum on a screen 1.00 m from the slit?


1
Expert's answer
2021-02-25T17:43:28-0500

Let's write the condition for the minimum (or destructive interference):


"wsin\theta_1=m\lambda,"

here, "w" is the width of the slit, "\theta_1" is the half-angular width of the central maximum, "\lambda" is the wavelength.

For "m=1" (the first minimum) our formula becomes:


"wsin\theta_1=\lambda,""sin\theta_1=\dfrac{\lambda}{w}=\dfrac{632.8\cdot10^{-9}\ m}{0.3\cdot10^{-3}\ m}=2.11\cdot10^{-3}."

Then, we can find the position of the minimum on the screen from the geometry:


"tan\theta_1\approx sin\theta_1\approx \theta_1\approx \dfrac{y}{L},""y=Lsin\theta_1=1\ m\cdot2.11\cdot10^{-3}=2.11\cdot10^{-3}\ m."

Finally, we can find the width of the central maximum:


"w_{central}=2y=2\cdot2.11\cdot10^{-3}\ m=4.22\cdot10^{-3}\ m=4.22\ mm."